OpenStudy (anonymous):

Maya plants 10 flower bulbs in the fall: 2 yellow,4 purple,and 4 pink. Find the probability that the first flower to bloom is pink

OpenStudy (anonymous):

this is n0. pinks / total no. of flowers = 4/10 = 2/5

OpenStudy (anonymous):

Given this information, and that each flower has the same probability of blooming in the fall, we can basically look at this as the probability of picking one of the 10 flowers out of a bag

OpenStudy (anonymous):

Jimmyrep gave the idea, and the answer